OpenStudy (anonymous):

A model rocket blasts off and moves upward with an acceleration of 13m/s2 until it reaches a height of 28m , at which point its engine shuts off and it continues its flight in free fall. A) What is the maximum height attained by the rocket? B) What is the speed of the rocket just before it hits the ground? C) What is the total duration of the rocket's flight?

OpenStudy (jfraser):

Use the equations for kinematics

OpenStudy (jfraser):

Even though the rocket is in free fall once the engine pellets off, it's still moving upwards, but now it's experiencing only the acceleration of gravity, 9.8m/s2

OpenStudy (anonymous):

If theres more than one equation which one would I need to use??

OpenStudy (dan815):

y''=13 y'=13t+v, y'=13t as y'(0)=0 y=13/2t^2+c, y=13/2t^2 as y(0) = 0 (starting from the ground)

OpenStudy (dan815):

28=13/2 t^2 time = sqrtC(28*2/13) y'(time)=velocity @ that 28m point after accerleration shuts off

OpenStudy (dan815):

from there you have another parabolic motion problem now you have a inital upward velocity so your equation changes to

OpenStudy (dan815):

y=-9.8/2 * t^2 + y'(@time_time_From_earlier_calculation)*t+28 y'=0, solve for t, and then solve for y(that_time)=maximum height then for part b you have this problem,